# DOMAIN: milDly cOnservative Model-bAsed offlINe reinforcement learning

### v2

@@ Abstract
Model-based reinforcement learning (RL), which learns environment model from offline dataset and generates more out-of-distribution model data, has become an effective approach to the problem of distribution shift in offline RL. Due to the gap between the learned and actual environment, conservatism should be incorporated into the algorithm to balance accurate offline data and imprecise model data. The conservatism of current algorithms mostly relies on model uncertainty estimation. However, uncertainty estimation is unreliable and leads to poor performance in certain scenarios, and the previous methods ignore differences between the model data, which brings great conservatism. Therefore, this paper proposes a milDly cOnservative Model-bAsed offlINe RL algorithm (DOMAIN) without estimating model uncertainty to address the above issues. DOMAIN introduces adaptive sampling distribution of model samples, which can adaptively adjust the model data penalty. In this paper, we theoretically demonstrate that the Q value learned by the DOMAIN outside the region is a lower bound of the true Q value, the DOMAIN is less conservative than previous model-based offline RL algorithms and has the guarantee of security policy improvement. The results of extensive experiments show that DOMAIN outperforms prior RL algorithms on the D4RL dataset benchmark, and achieves better performance than other RL algorithms on tasks that require generalization.
@@ I Introduction
Reinforcement learning (RL) aims to maximize user-defined rewards and acquire optimal behavioral skill. Since the 1970s, RL research has attracted widespread attention, and the advent of deep neural networks has greatly facilitated its development. Today, RL has been applied in various fields, including robot control [1, 2], autonomous driving [3], recommendation systems [4], and resource scheduling [5].
The paradigm of online RL involves an agent interacting with environment directly through trial-and-error[6]. However, online RL suffers from low sampling efficiency and high costs, making it impractical in certain scenarios [7]. For instance, in situations such as surgery and autonomous driving, real-time interaction between the agent and the environment may pose a threat to human safety. Since the robot may damage its components or surrounding objects, the cost of trial-and-error can be prohibitive in robot control [8]. Offline RL can effectively address these issues by learning policy from a static dataset without direct interaction with the environment.
Offline RL faces a significant challenge due to distribution shift, which is caused by differences between the policies in offline dataset and those learned by the agent. The selection of out-of-distribution (OOD) actions introduces extrapolation errors that may lead to unstable training of the agent [9]. To address the above issue, two types of methods have been developed: policy constraints [10, 11, 12] and value penalization [13, 14]. Huang et al. [40] proposed a mild policy evaluation method by constraining the difference between the $Q$ values of actions supported by target policy and those of actions contained within offline dataset. Yang et al. [41] treated different samples with different policy constraint intensities, which effectively improves the agent performance. The goal of the above methods is to minimize the discrepancy between the learned policy and that in the dataset. While they achieve remarkable performance, these methods fail to consider the effects of OOD actions, hampering the exploration and performance improvement. In contrast, Lyu et al. [15] proved that the agent performance can be effectively enhanced by exploring OOD region.
The aforementioned offline RL methods are model-free, which constrains exploration in OOD region, hindering performance improvements for agents. In contrast, model-based offline RL enables agents to continually interact with a trained environment model, generating data with broader coverage, thereby enhancing exploration in OOD region [16]. Moreover, agents acquire imaginative abilities through environmental modeling, enabling effective planning and improving control efficacy [17, 18]. Since the uncertainty of environment model, exploration by agents also carries risks. It is critical to strike a balance between return and risk.
Performing conservative policy optimization has become an effective way to the above problem. To incorporate the conservatism into policy, the typical method is estimating model uncertainty [19, 20, 39] or penalizing state-action pairs with high uncertainty [24, 25] during training. Experimental results have shown that model-based offline RL methods outperform most current model-free offline RL methods. However, measuring model uncertainty through neural networks results in prediction errors and low reliability. To address this issue, Yu et al. [21] proposed the COMBO algorithm based on CQL [13], which regularizes model data to suppress unreliable data and enhance reliable offline data. Rigter et al. [24] and Bhardwaj et al. [25] used an adversarial training framework, forcing the agent to act conservatively in OOD region.
@@
However, the above methods ignore differences between the model data and are overly conservative. Therefore, we design an adaptive sampling distribution of model data, and propose a mildly conservative model-based offline RL algorithm (DOMAIN) without estimating model uncertainty (as illustrated in Fig. 1). Here, the mildly conservatism is reflected in the following: higher exploration in OOD region and better performance of the agent. This algorithm imposes heavier penalties on model data with high errors while reducing the penalty for accurate model data to alleviate conservatism. Our main contributions are summarized as follows:
@@
    A new model-based offline RL method is proposed that adaptively adjusts the model data penalty, named DOMAIN, where the adaptive sampling distribution of model samples is designed. To our best knowledge, this is the first to incorporate adaptive sampling distribution to conservative value estimation in model-based offline RL.
@@
    Theoretical analyses demonstrate that the value function learned by the DOMAIN in the OOD region is a lower bound of the true value function, the DOMAIN is less conservative than previous offline model-based RL algorithms, and DOMAIN has a safety policy improvement guarantee.
@@
    Extensive experiments indicate that DOMAIN outperforms prior RL algorithms on the D4RL dataset benchmark, and achieves the best performance than other RL algorithms on tasks that require generalization.
The framework of this paper is as follows: Section 2 introduces the fundamental theory of RL. Section 3 shows the framework and implementation details of the DOMAIN algorithm. Section 4 gives some relevant theoretical analysis. Section 5 presents the algorithm performance on D4RL and generalization tasks. Finally, Section 6 summarizes the entire work and provides directions for future research.
@@ II-C Model-based Offline RL
However, offline dataset fails to cover the entire state-action space, resulting in disparities between the learned environment model $\widehat{T}_{\phi}\left(s^{\prime},r\mid s,a\right)$ and the actual environment model $T\left(s^{\prime},r\mid s,a\right)$, which leads to prediction errors in the model. To address these challenges, Janner et al. [23] employed a model ensemble approach to reduce biases introduced by probabilistic networks. Yu et al. [19] proposed the MOPO algorithm, which used model uncertainty as reward penalizes. The MOReL algorithm [20], similar to MOPO, incorporated penalty terms in rewards for OOD region data. Chen et al. [39] proposed offline model-based adaptable policy learning method that can adapt its behavior in out-of-support regions instead of learning in in-support regions. Furthermore, Yu et al. [21] considered the estimation errors of neural networks for model uncertainty and proposed the COMBO algorithm. This method utilizes value function regularization, aiming to push down unreliable model data and push up reliable offline data.
@@ III Methodology
Incorporating model uncertainty as a penalty in reward can alleviate the impact of imprecise model data. However, accurately estimating model uncertainty remains a challenging task. The COMBO algorithm effectively achieves a balance between return and risk by incorporating regularization based on CQL without estimating model uncertainty. Nonetheless, prevailing methodologies ignore the correlation between the magnitude of model errors and the sampling distribution. Consequently, such approaches tend to adopt overly conservative policies. Therefore, based on COMBO, a mildly conservative algorithm for model-based offline RL (DOMAIN) is proposed in this study.
@@ III-A Mildly Conservative Value Estimation
where $\lambda$ is a scaling factor used to adjust the weight of the regularization term, $\rho(s,a)$ denotes the sampling distribution of the model data, and $\mathcal{H}(\rho)$ is a constraint term for $\rho(s,a)$. The term $\mathbb{E}_{s,a\sim\rho(s,a)}[Q(s,a)]$ is employed to penalize unreliable model data, while $\mathbb{E}_{s,a\sim\mathcal{D}_{\text{offline }}}[Q(s,a)]$ is used to reward reliable offline data. The term $\mathcal{L}_{\beta}(Q,\widehat{\mathcal{T}}^{\pi}\widehat{Q})$ represents the Bellman error. In the Actor-Critic framework, the goal of policy evaluation is to minimize the Bellman error [26]. In the DOMAIN algorithm, the Bellman error consists of two components: the Bellman error from the offline data and the model data. It can be expressed as follows:
@@
where $f$ can be determined by two guidelines. One is the error of model data, and the other is the algorithm itself. If the model data is accurate and the algorithm can effectively penalize model data with large errors, $f$ can be set larger. On the contrary, if the error of model data is large and the algorithm fails to effectively penalize them, $f$ needs to be set smaller to guarantee the stability and security of the agent.
The value function $Q$ and the policy $\pi$ in Eq. (4) can be approximated using neural networks, denoted as $Q_{\omega}$ and $\pi_{\theta}$, respectively. $\widehat{\mathcal{T}}^{\pi}$ represents the actual Bellman operator used in the algorithm, while $\mathcal{T}_{\overline{\mathcal{M}}}^{\pi}$ and $\mathcal{T}_{\widehat{\mathcal{M}}}^{\pi}$ represent the operator applied under the empirical MDP and the learned model, respectively. The value function update follows the SAC framework [27], and thus the Bellman operator can be further expressed as:
@@
By optimizing Eq. (9), a mildly conservative value estimation can be achieved, and if the sampling distribution of model data $\omega(s,a)$ is known, the above expression can be optimized directly. Moreover, $\omega(s,a)$ reflects the degree of penalization for model data. Note that mildly conservative value estimation for model data doesn’t mean a more accurate value estimation. Since the errors in model data, the agent may explore dangerous areas. The penalty is ofthen added to the $Q$ value of the inaccurate model data to alleviate this problem, which causes that the estimated value is lower than the true value for inaccurate model data. The following part provides further derivation for model data sampling distribution.
@@ III-B Adaptive Value Regularization on Model Data
where $T_{\widehat{\mathcal{M}}}$ represents the state transition model in the learned environment $\widehat{\mathcal{M}}$, while $T_{\mathcal{M}}$ corresponds to that in the true environment $\mathcal{M}$. Since sampling distribution for model data needs to satisfy the equation: $\int_{\mathcal{S}}\int_{\mathcal{A}}\omega(s,a)\mathrm{d}s\mathrm{d}a=1$, the sampling distribution $\omega(s,a)$ based on finite model dataset can be denoted as:
${\color[rgb]{0,0,0}\omega(s,a)=\frac{g(s,a)}{\int_{\mathcal{S}}\int_{\mathcal{A}}g\left(s^{\prime},a^{\prime}\right)\mathrm{d}s^{\prime}\mathrm{d}a^{\prime}}}$ (11)
@@
As shown in the above equation, it is easy to find the probabilities of the data pair $\left(s,a,s^{\prime}\right)$ and $\left(s,a\right)$ being model data and offline data, respectively, are the key factors. To estimate these probabilities, two discriminators are constructed, namely $D_{\Phi_{sas}}\left(\cdot\mid s,a,s^{\prime}\right)$ and $D_{\Phi_{sa}}\left(\cdot\mid s,a\right)$, which approximate the probabilities using parameters $\Phi_{sas}$ and $\Phi_{sa}$, respectively. The network loss function employed for training the discriminators is the cross-entropy loss [30]:
@@
By training the discriminator networks, the gap between the learned and the actual environment can be estimated, leading to the estimation of the sampling distribution. Based on this, the different sampling probabilities among model data lead to adaptive value regularization for the model data.
@@ III-C Algorithm Details of DOMAIN
To incorporate conservatism into algorithm, the typical method is applying reward penalties for OOD data by uncertainty quantification directly [19, 20] or value function regularization indirectly [24, 25]. However, the above methods ignore differences between the model data, which may prevent the agent from choosing a better action since the $Q$ value is over-conservative. DOMAIN imposes heavier penalties on model data with high errors while reducing the penalty for accurate model data to alleviate conservatism. By reducing the penalty on the $Q$ value of model data with small errors, the agent may choose action $a$ with higher reward $r$ through $\arg\max_{a}Q(s,a)$, thus improving the performance of the learned policy.
Our algorithm DOMAIN, summarized in Algorithm 1, is based on SAC[27] and MOPO[19]. The training approach for the environment model closely resembles that of MOPO, employing a maximum likelihood estimation approach. $N$ dynamic environment models are trained and denoted as $\{\widehat{T}_{\phi}^{i}=\mathcal{N}[\mu_{\phi}^{i}(s,a),\sum_{\phi}^{i}(s,a)]\}_{i=1}^{N}$. Subsequently, a subset of $M$ models with superior training accuracy is selected, and their predicted values are aggregated to yield the final prediction. Eq.s (5) and (6) adhere to the well-established SAC framework, with the parameter $\alpha$ being automatically adapted throughout the training procedure [27]. The weight $\lambda$ of the regularization term and the rolling horizon $H$ exert a substantial influence on the agent training.
It employs Eq. (9) to perform mildly conservative policy evaluation, Eq. (6) for policy improvement, and Eq. (13) to train the discriminator. Here, $g(s,a)=\mathbb{E}_{s^{\prime}\sim\widehat{T}_{\widehat{\mathcal{M}}}}\left[T_{\widehat{\mathcal{M}}}\left(s^{\prime}\mid s,a\right)/T_{\mathcal{M}}\left(s^{\prime}\mid s,a\right)\right]$ represents the expectation of the dynamic ratio. This expectation is estimated by sampling $m$ states from a Gaussian distribution $\widehat{T}_{\phi}\left(s^{\prime},r\mid s,a\right)=\mathcal{N}[\mu_{\phi}(s,a),\sum_{\phi}(s,a)]$ obtained from the trained model. The computation of sampling distribution $\omega(s,a)$ for model data is based on the finite model dataset $\mathcal{D}_{\text{model}}$, which fails to cover all action space. Therefore, in practical implementation, the estimated sampling distribution $\omega(s,a)$ is approximated by $g(s,a)/\sum_{(s^{\prime},a^{\prime})}g(s^{\prime},a^{\prime})$.
Note that the above training of algorithm DOMAIN only depends on the offline dataset. If the offline dataset is a little, the agent performance is difficult to improve. The trained environment model is trained from the offline dataset. A small dataset will lead to a large gap between the true and estimated environments, which brings large model data errors. The proposed method will fail to achieve the balance between accurate offline data and inaccuracy model data. Therefore, the offline dataset cannot be too small and needs to cover a certain state-action space.
@@ IV Theoretical Analysis of DOMAIN
In this section, we analyze the DOMAIN algorithm and demonstrate that the adaptive sampling distribution of model data is equivalent to the adaptive adjustment of penalties for rewards. Furthermore, we prove that the learned value function in DOMAIN serves as a lower bound for the true value function in OOD region, effectively suppressing the value function in highly uncertain region. In comparison to COMBO, DOMAIN exhibits weaker conservatism. By employing a lower bound optimization strategy, DOMAIN guarantees safety policy improvement.
@@ IV-A Adaptive Adjustment of Penalties for Rewards
The detailed proof can be found in Appendix A. Theorem 1 shows that the operator $\widehat{\mathcal{T}}$ is a $\gamma$-contraction mapping in the $\mathcal{L}_{\infty}$ norm. According to the fixed point theorem, any initial $Q$ function can converge to a unique fixed point $Q^{*}$ by repeatedly applying the Bellman operator $\widehat{\mathcal{T}}^{\pi}$. By setting the derivative of the Eq. (9) concerning $Q^{k}$ to zero during the $k$-th iteration, the equation can be obtained:
@@
Here, the distribution for offline data sampling, denoted as $d(s,a)$, is commonly defined as $d(s,a)=d^{\pi_{b}}(s,a)=d^{\pi_{b}}(s)\pi_{b}(a|s)$, where $d^{\pi_{b}}(s,a)$ represents the marginal distribution of states and actions under the behavioral policy $\pi_{b}$. On the other hand, the data sampling distribution used for computing the Bellman error, denoted as $d_{\beta}(s,a)$, is defined as $d_{\beta}(s,a)=(1-f)d^{\pi_{b}}(s,a)+fd_{\widehat{\mathcal{M}}}^{\pi}(s,a)$. Here, $d_{\widehat{\mathcal{M}}}^{\pi}(s,a):=d_{\widehat{\mathcal{M}}}^{\pi}(s)\pi(a|s)$ represents the marginal distribution of states and actions under the learning policy $\pi$, and $d_{\widehat{\mathcal{M}}}^{\pi}(s)$ represents the discounted marginal state distribution when executing policy $\pi$ in the learned model $\widehat{\mathcal{M}}$.
Let us designate the second term in Eq. (14) as $\eta(s,a)$, which serves as the adjustment of penalties for rewards. When the model data $(s,a)$ exhibits high errors, the value of the designed sampling distribution $\omega(s,a)$ increases. This indicates that the data point $(s,a)$ either resides in the OOD region or a region with a relatively low density of offline dataset, implying that $d(s,a)$ approaches zero. Consequently, $\omega(s,a)>d(s,a)$, resulting in $\eta(s,a)>0$ that serves as a penalty term. Moreover, the magnitude of penalization in the Q value increases with larger errors in the model data. In contrast, when the model data error is small, suggesting that the data is situated in a region proximal to true environment, it is possible for $\omega(s,a)$ to be smaller than $d(s,a)$, leading to $\eta(s,a)<0$ which acts as a reward term. This encourages exploration in accurate model data region during training process.

### v3

@@ Abstract
Model-based reinforcement learning (RL), which learns environment model from offline dataset and generates more out-of-distribution model data, has become an effective approach to the problem of distribution shift in offline RL. Due to the gap between the learned and actual environment, conservatism should be incorporated into the algorithm to balance accurate offline data and imprecise model data. The conservatism of current algorithms mostly relies on model uncertainty estimation. However, uncertainty estimation is unreliable and leads to poor performance in certain scenarios, and the previous methods ignore differences between the model data, which brings great conservatism. Therefore, this paper proposes a milDly cOnservative Model-bAsed offlINe RL algorithm (DOMAIN) without estimating model uncertainty to address the above issues. DOMAIN introduces adaptive sampling distribution of model samples, which can adaptively adjust the model data penalty. In this paper, we theoretically demonstrate that the Q value learned by the DOMAIN outside the region is a lower bound of the true Q value, the DOMAIN is less conservative than previous model-based offline RL algorithms and has the guarantee of safety policy improvement. The results of extensive experiments show that DOMAIN outperforms prior RL algorithms on the D4RL dataset benchmark.
@@ I Introduction
Reinforcement learning (RL) has been widely applied in various fields, including robot control [1, 2], autonomous driving [3], recommendation systems [4], and resource scheduling [5]. The paradigm of online RL involves an agent interacting with environment directly through trial-and-error[6]. However, online RL suffers from low sampling efficiency and high interaction cost, making it impractical in certain scenarios [7]. Offline RL can effectively address the above problems by learning policy from offline dataset without interaction with the environment.
Offline RL faces a significant challenge due to distribution shift, which is caused by difference between the policy in offline dataset and those learned by the agent. The selection of out-of-distribution (OOD) data introduces extrapolation error that may lead to unstable training [9]. To address the above issue, two types of methods have been developed: policy constraints [10, 11, 12] and value penalization [13, 14]. The goal of the above methods is to minimize the discrepancy between the learned policy and that in the dataset. While they achieve remarkable performance, these methods are model-free and fail to consider the effects of OOD data, hampering the exploration and performance improvement.
Model-based offline RL, which enables agent to continually interact with a trained environment model, can broaden data coverage and enhance exploration in OOD region [16]. However, due to the gap between the estimated and true environment model, the generated data maybe inaccurate, that is the generated model data is inaccurate, which brings risks for agent training. The current methods often perform conservative policy optimization to solve this problem. The typical method is estimating model uncertainty [19, 20, 39] or penalizing state-action pairs with high uncertainty [24, 25]. However, estimating model uncertainty through neural networks brings prediction errors and low reliability. To address this problem, Yu et al. [21] proposed the COMBO algorithm based on CQL [13], which regularizes model data to suppress unreliable data and enhance reliable offline data.
@@
However, when regularizing model data, the current model-based offline RL methods fail to consider the differences between the model data and are overly conservative. For example, for precise model data, penalizing them as if they were imprecise model data can lead to underutilization of the precise data, thereby increasing the conservativeness of the algorithm. Therefore, we design an adaptive sampling distribution of model data, and propose a mildly conservative model-based offline RL algorithm (DOMAIN) without estimating model uncertainty (as illustrated in Fig. 1). Here, the mildly conservatism is reflected as: higher exploration in OOD region and better performance of agent. This algorithm imposes heavier penalties on model data with high errors while reducing the penalty for accurate model data. Our main contributions are summarized as follows:
@@
    A new model-based offline RL method, named DOMAIN, is proposed, where the adaptive sampling distribution of model samples is designed to adaptively adjusts the model data penalty.
@@
    Theoretical analysis demonstrates that the value function learned by the DOMAIN in the OOD region is a lower bound of the true value function, and DOMAIN has a safety policy improvement guarantee.
@@
    The results of extensive experiment indicate that DOMAIN can outperform previous RL algorithms on the D4RL dataset benchmark.
Notably, the exploration risk for unknown region can be reduced through underestimating the value of inaccurate model data. Through reducing the penalty on the $Q$ value of model data with small error, the agent can choose action $a$ with higher reward $r$ by $\arg\max_{a}Q(s,a)$, thus improving performance of the learned policy. The framework of this paper is as follows: Section 2 introduces the fundamental theory of RL. Section 3 shows the framework and implementation details of the DOMAIN algorithm. Section 4 gives some relevant theoretical analysis. Section 5 presents the algorithm performance on D4RL tasks. Finally, Section 6 summarizes the entire work.
@@ II-C Model-based Offline RL
However, offline dataset fails to cover the entire state-action space, resulting in disparities between the learned environment model $\widehat{T}_{\phi}\left(s^{\prime},r\mid s,a\right)$ and the actual environment model $T\left(s^{\prime},r\mid s,a\right)$. To address this challenge, Janner et al. [23] employed a model ensemble approach to reduce biases introduced by probabilistic networks. Yu et al. [19] proposed the MOPO algorithm, which used model uncertainty as reward penalizes. Yu et al. [21] considered the estimation errors of neural networks for model uncertainty and utilized value function regularization, aiming to push down unreliable model data and push up reliable offline data. However, the above methods fail to consider the difference between the model data and overly conservative, thus the algorithm DOMAIN is proposed in this paper.
@@ III Methodology
This section gives the details of mildly conservative model-based offline RL (DOMAIN) algorithm, including mildly conservative estimation for value function and adaptive value regularization on model data.
@@ III-A Mildly Conservative Value Estimation
where $\lambda$ is a scaling factor used to adjust the weight of the regularization term, $\rho(s,a)$ denotes the sampling distribution of the model data, and $\mathcal{H}(\rho)$ is a constraint term for $\rho(s,a)$. The term $\mathbb{E}_{s,a\sim\rho(s,a)}[Q(s,a)]$ is employed to penalize unreliable model data, while $\mathbb{E}_{s,a\sim\mathcal{D}_{\text{offline }}}[Q(s,a)]$ is used to reward reliable offline data. The term $\mathcal{L}_{\beta}(Q,\widehat{\mathcal{T}}^{\pi}\widehat{Q})$ represents the Bellman error.
In the Actor-Critic framework, the goal of policy evaluation is to minimize the Bellman error [26]. In the DOMAIN algorithm, the Bellman error, consisting of the Bellman error from the offline data and the model data, can be denoted as
@@
where $\widehat{\mathcal{T}}^{\pi}$ represents the actual Bellman operator used in the algorithm, while $\mathcal{T}_{\overline{\mathcal{M}}}^{\pi}$ and $\mathcal{T}_{\widehat{\mathcal{M}}}^{\pi}$ represent the operator applied under the empirical MDP and the learned model, respectively.
The value function $Q$ and the policy $\pi$ in Eq. (4) can be approximated using neural networks, denoted as $Q_{\omega}$ and $\pi_{\theta}$, respectively. The value function update follows the SAC framework [27], and thus the Bellman operator can be further expressed as:
@@
By optimizing Eq. (9), a mildly conservative value estimation can be achieved. The sampling distribution of model data $\omega(s,a)$ reflects the degree of penalization for model data. In below part, we give further derivation for the sampling distribution of model data $\omega(s,a)$.
@@ III-B Adaptive Value Regularization on Model Data
where $T_{\widehat{\mathcal{M}}}$ represents the state transition model in the learned environment $\widehat{\mathcal{M}}$, while $T_{\mathcal{M}}$ corresponds to that in the true environment $\mathcal{M}$. Since sampling distribution for model data needs to satisfy the equation: $\sum_{s,a}\omega(s,a)=1$, the sampling distribution $\omega(s,a)$ based on finite model dataset can be denoted as:
$\omega(s,a)={g(s,a)}\big{/}{\sum_{s^{\prime},a^{\prime}}g\left(s^{\prime},a^{\prime}\right)}$ (11)
@@
where $p\left(\text{model}|s,a,s^{\prime}\right)$ denotes the probability that the pair $\left(s,a,s^{\prime}\right)$ belongs to model data. To estimate the above probabilities, two discriminators are constructed, namely $D_{\Phi_{sas}}\left(\cdot\mid s,a,s^{\prime}\right)$ and $D_{\Phi_{sa}}\left(\cdot\mid s,a\right)$, which approximate the probabilities using parameters $\Phi_{sas}$ and $\Phi_{sa}$, respectively. The network loss function employed for training the discriminators is the cross-entropy loss [30]:
@@
By training the discriminator networks, the gap between the learned and the actual environment can be estimated. Then, the sampling distribution of model data can be obtained. Based on this, the different sampling probabilities among model data can bring the adaptive value regularization for the model data.
@@ III-C Algorithm Details of DOMAIN
DOMAIN imposes heavier penalty on model data with high errors while reducing the penalty for accurate model data to alleviate conservatism. By reducing the penalty on the $Q$ value of model data with small errors, the agent may choose action $a$ with higher reward $r$ through $\arg\max_{a}Q(s,a)$, thus improving the performance of the learned policy. The algorithm DOMAIN, summarized in Algorithm 1, is based on SAC [27] and MOPO [19]. $N$ dynamic environment models are trained and a subset of $M$ models with superior training accuracy is selected. Eq.s (5) and (6) adhere to SAC framework, with the parameter $\alpha$ being automatically adjusted [27].
DOMAIN employs Eq. (9) to perform mildly conservative policy evaluation, Eq. (6) for policy improvement, and Eq. (13) to train the discriminator. The term $g(s,a)$ represents the expectation of the dynamic ratio. This expectation is estimated by sampling $m$ states from a Gaussian distribution $\widehat{T}_{\phi}\left(s^{\prime},r\mid s,a\right)=\mathcal{N}[\mu_{\phi}(s,a),\sum_{\phi}(s,a)]$ obtained from the trained model. The weight $\lambda$ of the regularization term and the rolling horizon $H$ exert a substantial influence on the agent training. The detailed analysis can be found in section V-D.
@@ IV Theoretical Analysis of DOMAIN
In this section, we demonstrate that the adaptive sampling distribution of model data is equivalent to the adaptive adjustment of penalties for rewards. Furthermore, we prove that the learned value function in DOMAIN serves as a lower bound for the true value function in OOD region, effectively suppressing the value function in highly uncertain region. By employing a lower bound optimization strategy, DOMAIN guarantees safety policy improvement.
@@ IV-A Adaptive Adjustment of Penalties for Rewards
The detailed proof can be found in Appendix A. Theorem 1 shows that the operator $\widehat{\mathcal{T}}$ is a $\gamma$-contraction mapping in the $\mathcal{L}_{\infty}$ norm. According to the fixed point theorem, any initial $Q$ function can converge to a unique fixed point $Q^{*}$ by repeatedly applying the Bellman operator $\widehat{\mathcal{T}}^{\pi}$.
By setting the derivative of the Eq. (9) concerning $Q^{k}$ to zero during the $k$-th iteration, the following equation can be obtained:
@@
where the distribution for offline data sampling, denoted as $d(s,a)$, is defined as $d(s,a)=d^{\pi_{b}}(s,a)=d^{\pi_{b}}(s)\pi_{b}(a|s)$, and $d^{\pi_{b}}(s,a)$ represents the marginal distribution of states and actions under the behavioral policy $\pi_{b}$. The data sampling distribution used for computing the Bellman error, denoted as $d_{\beta}(s,a)$, is defined as $d_{\beta}(s,a)=(1-f)d^{\pi_{b}}(s,a)+fd_{\widehat{\mathcal{M}}}^{\pi}(s,a)$. Here, $d_{\widehat{\mathcal{M}}}^{\pi}(s,a):=d_{\widehat{\mathcal{M}}}^{\pi}(s)\pi(a|s)$ represents the marginal distribution of states and actions under the learning policy $\pi$, and $d_{\widehat{\mathcal{M}}}^{\pi}(s)$ represents the discounted marginal state distribution when executing policy $\pi$ in the learned model $\widehat{\mathcal{M}}$.
We denote the second term in Eq. (14) as $\eta(s,a)$, which serves as the adjustment of penalties for rewards. When the model data $(s,a)$ exhibits high errors, the value of the designed sampling distribution $\omega(s,a)$ increases. This indicates that the data point $(s,a)$ either resides in the OOD region or a region with a relatively low density of offline dataset, implying that $d(s,a)$ approaches zero. Consequently, $\omega(s,a)>d(s,a)$, resulting in $\eta(s,a)>0$ that serves as a penalty term. Moreover, the magnitude of penalization in the Q value increases with larger errors in the model data. In contrast, when the model data error is small, that is the data is situated in a region proximal to true environment, it is possible for $\omega(s,a)$ to be smaller than $d(s,a)$, leading to $\eta(s,a)<0$ which acts as a reward term. This encourages exploration in accurate model data region during training process.